Shanon and Piper are on the golf team. Shanon buys 10 cheap golf balls and 7 expensive ones for $31.75. Piper buys 8 cheap and 5 expensive balls for $23.75. Find the prices of the two kinds of golf balls.
Answer by mananth(16946)

cheap ball be $x
expensive ball be $y
Shanon - cost equation
10x+7y= 31.75
Piper - cost equation
8x+5y= 23.75
8(10x+7y)-10(8x+5y)= 8*31.75 - 10*23.75
80x+56y-80x-50y=252-237.5
6y=14.5
y=2.42
Plug the value of in one equation
10x+7*2.42 =31.75
10x= 31.75-16.94
10x= 14.81
x=1.48
CHECK
10*1.48+7*2.42=31.74